# GraphBench: Next-generation graph learning benchmarking

### v1

@@ 1 Introduction
Machine learning on graphs using graph neural networks (GNNs), specifically message-passing neural networks (MPNNs) (Gil+2017; Sca+2009) and graph transformers (GTs) (Mue+2023), has become a cornerstone of modern machine learning research, spanning applications in drug design (Won+2023), recommender systems (Yin+2018a), chip design (deepgate2025), and combinatorial optimization (Cap+2021). Despite these advances, benchmarking in the field remains highly fragmented. Narrow, task-specific datasets and inconsistent evaluation protocols limit reproducibility and hinder meaningful comparison across tasks and domains (Spe+2025). As Spe+2025 argue, current benchmarks tend to emphasize specific domains—such as 2D molecular graphs—while overlooking more impactful, real-world applications like relational databases, chip design, and combinatorial optimization. Many benchmark datasets fail to accurately reflect the complexity of real-world structures, resulting in inadequate abstractions and misaligned use cases. In addition, the field’s heavy reliance on accuracy as a primary metric encourages overfitting, rather than fostering generalizable, robust models. As a result, these limitations have significantly hindered progress in graph learning.
@@
We introduce GraphBench, a next-generation benchmarking suite addressing key shortcomings in graph learning evaluation. GraphBench spans diverse domains—algorithmic reasoning, chip design, combinatorial optimization, SAT solver prediction, social networks, and weather forecasting—and supports node-, edge-, graph-level, and generative tasks. It provides standardized splits, domain-specific metrics, and hyperparameter-tuning scripts, along with OOD evaluation for selected datasets to assess generalization. We benchmark modern MPNNs and graph transformers, offering strong baselines and insights into architectural trade-offs. By unifying heterogeneous tasks under one framework, GraphBench enables reproducible, robust, and impactful graph learning research, see Figure˜1, paving the way for graph foundation models.
@@ 2.1 Example usage of GraphBench
We provide a unified and easy-to-use interface for GraphBench datasets. Inspired by benchmarks such as OGB (hu2020ogb) and LRGB (dwivedi2022long), we use wrapper methods to handle dataset loading. Unlike competing benchmarks, however, GraphBench offers a single loading class (graphbench.Loader) for all datasets, independent of task type. Each wrapper includes the necessary precomputations and comes with predefined splits, enabling direct use in downstream tasks. Nonetheless, we offer customization options for each dataset integrated in the loader method. For seamless integration, all datasets are fully compatible with PyTorch Geometric through its InMemoryDataset interface. We provide GraphBench as a python package distributed via PyPi under the name graphbench-lib, allowing us to provide continuous updates to code and datasets.
@@ 3.1.1 Social networks: Predicting engagements on BlueSky
Social networks—graphs with users as nodes and interactions as edges—provide a structurally rich domain for graph-based machine learning (newman2010networks). They exhibit hubs, peripheral nodes, clustered communities, and evolving, often directed edges. Dynamic user interests create temporal dependencies that shape representations and predictions. These properties make social networks valuable graph learning benchmarks, requiring models, e.g., to handle edge directionality and structural diversity. In particular, our proposed benchmark involves the prediction of the number of engagements a user will receive on upcoming posts, i.e., a user-centric measure of near-term influence, with applications in ranking users, trend detection, influencer marketing, and proactive moderation.
@@
Given timestamps $t_{0}<t_{1}<t_{2}$, user $u$ and target engagement $\kappa$, the learning task is to predict value $y^{\kappa\ziffer@CommaOri(u)}_{\tau_{1\ziffer@CommaMy 2}}$ given the overall interaction graph $G^{\iota}_{\tau_{0\ziffer@CommaMy 1}}$ and post contents $\{T^{(u)}_{\tau_{0\ziffer@CommaMy 1}}\mid u\in V\}$. We assess model performance using the mean absolute error (MAE), coefficient of determination $R^{2}$, and the Spearman correlation $\sigma$, where the last two are calculated specifically for each target engagement $\kappa$. We note that $\sigma$ is relevant to us since it measures how well predictions preserve the relative ordering of users by future engagements, a key criterion when ranking influence in heavy-tailed social networks, not necessarily captured by MAE and $R^{2}$. See more on the evaluation metrics in Section˜B.1.
@@
    While ensuring the model is not trained on post contents and network interactions from the future, by setting $t_{0}=t_{\text{start}}$ in all splits this partitioning strategy also reflects a realistic scenario where a social network “grows in time,” in the sense that user representations evolve as they generate new content and their connections expand as they interact with more users. See Section˜B.1 for details on the exact values of $t_{\text{start}}\ziffer@CommaOri t_{\text{end}}\ziffer@CommaOri t_{0}\ziffer@CommaOri\ldots\ziffer@CommaOri t_{2}$.
@@
We assess model performance on three different interaction types: quotes, replies, reposts and for three different engagement types: likes, replies, and reposts. As for baselines in comparison we included DeepSets (zaheer2017deep), MLPs, and MPNNs. Note how the first two are connectivity-agnostic; differently than MLPs, DeepSets access, however, global feature information, see more in Section˜C.3.
@@ Results
Table˜3 shows a clear, consistent ordering from DeepSets to MLPs to MPNNs, across all interaction types in terms of lower MAE, higher $R^{2}$, and Spearman ($\sigma$). These results suggest that local, directed relational structures are informative for predicting engagements over a short time horizon, as both DeepSets and MLPs, which ignore graph structure, lag behind MPNNs. However, it has to be noted the global feature information accessed by DeepSets appears detrimental. We also note that the absolute values of $R^{2}$ and $\sigma$ remain relatively low, indicating that the proposed tasks are far from being solved and leaving room for meaningful gains through more advanced, task-specific MPNNs.
@@ 3.2.1 Chip design: Learning to generate small circuits
Recently, machine learning has begun to influence key stages of chip design, including macro placement (alphachip) and floor planning (mallappa2024floorset). In this work, we propose focusing such efforts on logic synthesis, framing it as a conditional graph generation problem, i.e., generating a circuit that correctly implements a given Boolean function while minimizing the number of gates. The goal is thus to produce circuits that satisfy a prescribed truth table exactly while optimizing structural efficiency.
@@
Formally, let $f\colon\{0\ziffer@CommaMy 1\}^{n}\to\{0\ziffer@CommaMy 1\}^{m}$ be a Boolean function with $n$ inputs and $m$ outputs. Its behavior is fully specified by a truth table $\mathbold{T}^{f}\in\{0\ziffer@CommaMy 1\}^{m\times 2^{n}}$, where each column $\mathbold{T}^{f}_{\mathbold{x}}$ corresponds to an input $\mathbold{x}\in\{0\ziffer@CommaMy 1\}^{n}$ and satisfies $f(\mathbold{x})=\mathbold{T}^{f}_{\mathbold{x}}$. A logic circuit $C\coloneqq(V\ziffer@CommaOri E)$ is represented as a directed acyclic graph (DAG) of in-degree at most two. The node set is partitioned into input nodes $V^{\text{in}}\coloneqq\{v^{\text{in}}_{1}\ziffer@CommaOri\ldots\ziffer@CommaOri v^{\text{in}}_{n}\}$, gate nodes $V^{\text{gate}}\coloneqq\{v^{\text{gate}}_{1}\ziffer@CommaOri\ldots\ziffer@CommaOri v^{\text{gate}}_{k}\}$, each computing a Boolean function of arity at most two (e.g., And), and output nodes $V^{\text{out}}\coloneqq\{v^{\text{out}}_{1}\ziffer@CommaOri\ldots\ziffer@CommaOri v^{\text{out}}_{m}\}$, each of in-degree one. The edge set $E\subseteq V\times V\times\{0\ziffer@CommaMy 1\}$ specifies connections, where a label of $1$ denotes signal inversion (Not) and 0 denotes direct transmission. For an input $\mathbold{x}\in\{0\ziffer@CommaMy 1\}^{n}$, values are assigned to input nodes, propagated through gate nodes (applying inversion where specified), and collected at output nodes to yield the circuit output $C(\mathbold{x})\in\{0\ziffer@CommaMy 1\}^{m}$. Because the combination of And and Not is functionally complete, any Boolean function can be represented using only these two operations. We therefore adopt the standard representation of and-inverter graphs (AIGs), where circuits are DAGs of And gates connected by edges that may invert their signals (mishchenko2005fraigs). AIGs are widely used in industrial logic synthesis and provide a compact, canonical representation of circuits.
@@ 3.2.2 Electronic circuits: Predicting voltage conversion ratio
Analog circuit design remains critical but labor-intensive in EDA (Zhang2019CircuitGNNGN; Zhang2020GRANNITEGN; zhao2020automated). Unlike mature digital flows, analog design still relies on expert intuition due to sensitivity to device variations and complex constraints. This is especially true for custom power-converter design, where topology generation and parameter tuning must balance conflicting metrics. Even with a modest number of components, hundreds of thousands of topologies are possible; yet, new converter families are mostly handcrafted. Automating this process could accelerate design cycles and expand the exploration of novel architectures. Learning-based methods offer promise as surrogate models that predict performance (Fan2021FromST; fan2024graph), enabling faster search over vast spaces. Power converters’ natural graph structure makes them an ideal benchmark, highlighting robustness to structural sensitivity and the ability to cope with extreme class imbalance.
@@ 3.3.1 SAT solving: Algorithm selection and performance prediction
In GraphBench, we introduce, to the best of our knowledge, the largest algorithm-selection benchmark for SAT solvers, covering eleven solvers and more than 100 000 instances. Our benchmark supports multiple graph representations of SAT formulae and also provides the SATzilla 2024 feature set (ShaHoo24), enabling their combination with MPNNs. Unlike existing GNN benchmarks for SAT (LiEtAl24), which primarily focus on satisfiability prediction or assignment generation, our benchmark is designed for practical downstream tasks in SAT solving—specifically, algorithm selection and performance prediction.
@@
We created the largest algorithm selection scenario for SAT solvers available to date. We collected all SAT Competition instances via GBD (IseJab24) and added all instances from the AClib benchmark for algorithm configuration (HutEtAl14b). Together, these sources include SAT formulae from diverse real-world applications and synthetically generated instances. To further enrich the dataset, we generated additional formulae using three well-established SAT instance generators, i.e.,
@@
We provide three datasets, small, which includes only small formulae with up to 3 000 variables and 15 000 clauses, medium, which consists of the small formulae with additional medium size formulae with size of up to 20 000 variables and 80 000 clauses and large, which includes all formulae. This way, we provide datasets with different hardness levels, as the size of the formula can cause hurdles related to training time and GPU memory. While the large and medium datasets are unlikely to be accessible to current hardware, we provide it to allow the benchmark to stay relevant in the long term.
@@
All SAT solving tasks in the benchmark are graph-level. For the graph transformer baselines, we employ an encoder-processor-decoder pipeline, as described in Appendix˜C, with no edge or node encoding. While we considered GT for the tasks, we encountered high computational costs for calculating the positional embedding (more than a CPU day) and therefore excluded it from the evaluation. Furthermore, when training on medium- and large-sized graphs, we encountered out-of-memory errors even on an H100 with mixed-precision training. Consequently, we only consider small formulae for MPNNs. Similarly, we observed out-of-memory with the clause graphs, and therefore excluded them from the evaluation.
@@ Results
We start by presenting the results for the performance prediction task. We present the RMSE of the log-scaled computation time for the solvers from SAT Competition 2024 (Kissat, BreakIDKissat, KissatMABDC, AMSAT) for the small formulae in table Table˜8. The table shows that traditional empirical performance predictors (random forest and XGBoost) that use hand-crafted features outperform MPNNs by a high margin. Furthermore, we observe significant differences between the various graph types. While usage of LCG is common as an input for MPNNs in the context of SAT solving (SelEtAl19; TonGro25), the smaller VCG and VG perform better, which shows that the choice of input graph can have a high influence on the results. We furthermore present the results of random forest and XGBoost on medium-size and large formulae in Tables˜9 and 9, such that they can be compared against graph machine learning solutions in the future. Overall, we observe that the random forest is the best-performing EPM for SAT solvers. For the algorithm selection task, we report the closed gap metric of different selectors in Table˜10. The VG representation allows for better performance than Both VCG and LCG, although using GIN across all three graph representation approaches, underperform pairwise regression with hand-crafted features. Furthermore, VG also outperforms multi-class classification and achieves a slightly smaller gap closure than pairwise classification. In contrast, using LCG leads to worse performance than the single best solver, indicating poor generalization to unseen instances. In their work, ZhaEtAl24 claimed that their MPNN-based method outperforms traditional approaches. However, there are several differences: the instance and solver composition differ, they used a SAT-specific MPNN layer, whereas we employ a general-purpose GIN, and our baselines use the latest SATzilla 2024 features, which have shown significantly better results than the previous version (ShaHoo24). Since the authors did not publish their code, we were unable to include their method as a baseline in our study. Among the three baselines that are based on hand-crafted features, we observe that multi-class classification underperforms both pairwise classification and regression, reaching a negative closed gap on the small and medium-sized datasets. This highlights that algorithm selection is not a simple multi-class classification problem. In that formulation, the model loses valuable information about the performance of algorithms other than the best one. Finally, we find more gap closed for the large dataset across all three baselines, as it includes harder instances. As times increase, the benefit of algorithm selection grows, since feature computation accounts for a smaller fraction of total computation time. Moreover, for harder instances, multiple solvers frequently time out, so choosing a solver that successfully solves the instance becomes especially valuable, given the additional penalty imposed by the PAR score.

### v2

@@ 1 Introduction
Machine learning on graphs using graph neural networks (GNNs), specifically message-passing neural networks (MPNNs) (Gil+2017; Sca+2009) and graph transformers (GTs) (Mue+2023), has become a cornerstone of modern machine learning research, spanning applications in drug design (Won+2023), recommender systems (Yin+2018a), chip design (zheng2025deepgate), and combinatorial optimization (Cap+2021). Despite these advances, benchmarking in the field remains highly fragmented. Narrow, task-specific datasets and inconsistent evaluation protocols limit reproducibility and hinder meaningful comparison across tasks and domains (Spe+2025). As Spe+2025 argue, current benchmarks tend to emphasize specific domains—such as 2D molecular graphs—while overlooking more impactful, real-world applications like relational databases, chip design, and combinatorial optimization. Many benchmark datasets fail to accurately reflect the complexity of real-world structures, resulting in inadequate abstractions and misaligned use cases. In addition, the field’s heavy reliance on accuracy as a primary metric encourages overfitting, rather than fostering generalizable, robust models. As a result, these limitations have significantly hindered progress in graph learning.
@@
We introduce GraphBench, a next-generation benchmarking suite addressing key shortcomings in graph learning evaluation. GraphBench spans diverse domains—algorithmic reasoning, chip design, combinatorial optimization, SAT solver prediction, social networks, and weather forecasting—and supports node-, edge-, graph-level, and generative tasks. It provides standardized splits, domain-specific metrics, and hyperparameter-tuning scripts, along with out-of-distribution evaluation for selected datasets to assess generalization. We benchmark modern MPNNs and graph transformers, offering strong baselines and insights into architectural trade-offs. By unifying heterogeneous tasks under one framework, GraphBench enables reproducible, robust, and impactful graph learning research, see Figure˜1, paving the way for graph foundation models.
@@ 2.1 Example usage of GraphBench
We provide a unified and easy-to-use interface for GraphBench datasets. Inspired by benchmarks such as OGB (hu2020ogb) and LRGB (dwivedi2022long), we use wrapper methods to handle dataset loading. Unlike competing benchmarks, however, GraphBench offers a single loading class (graphbench.Loader) for all datasets, independent of task type. Each wrapper includes the necessary precomputations and comes with predefined splits, enabling direct use in downstream tasks. Nonetheless, we offer customization options for each dataset integrated in the loader method. For seamless integration, all datasets are fully compatible with PyTorch Geometric through its InMemoryDataset interface. We provide GraphBench as a Python package distributed via PyPi under the name graphbench-lib, enabling continuous updates to code and datasets.
@@ 3.1 Social sciences
Figure 2: Excerpt of a dynamic social network over three timesteps $t_{a}<t_{b}<t_{c}$ constructed from the BlueSky dataset. Node features are displayed by color gradient, and the target value is encoded by node size. Edges may be rewired between timesteps.
@@ 3.1.1 Social networks: Predicting engagements on BlueSky
Social networks—graphs with users as nodes and interactions as edges—provide a structurally rich domain for graph-based machine learning (newman2010networks). They exhibit hubs, peripheral nodes, clustered communities, and evolving, often directed edges. Dynamic user interests create temporal dependencies that shape representations and predictions. These properties make social networks valuable benchmarks for graph learning, requiring models to handle edge directionality and structural diversity. In particular, our proposed benchmark involves predicting the number of engagements a user will receive on upcoming posts, i.e., a user-centric measure of near-term influence, with applications in user ranking, trend detection, influencer marketing, and proactive moderation. Figure˜2 provides an overview of how the graph is constructed from the BlueSky dataset and changes over time.
@@
Given timestamps $t_{0}<t_{1}<t_{2}$, user $u$ and target engagement $\kappa$, the learning task is to predict value $y^{\kappa\ziffer@CommaOri(u)}_{\tau_{1\ziffer@CommaMy 2}}$ given the overall interaction graph $G^{\iota}_{\tau_{0\ziffer@CommaMy 1}}$ and post contents $\{T^{(u)}_{\tau_{0\ziffer@CommaMy 1}}\mid u\in V\}$. We assess model performance using the mean absolute error (MAE), coefficient of determination $R^{2}$, and the Spearman correlation $\sigma$, where the last two are calculated specifically for each target engagement $\kappa$. We note that $\sigma$ is relevant to us since it measures how well predictions preserve the relative ordering of users by future engagements, a key criterion when ranking influence in heavy-tailed social networks, not necessarily captured by MAE and $R^{2}$. See more on the evaluation metrics in Section˜A.1.
@@
    While ensuring the model is not trained on post contents and network interactions from the future, by setting $t_{0}=t_{\text{start}}$ in all splits this partitioning strategy also reflects a realistic scenario where a social network “grows in time,” in the sense that user representations evolve as they generate new content and their connections expand as they interact with more users. See Section˜A.1 for details on the exact values of $t_{\text{start}}\ziffer@CommaOri t_{\text{end}}\ziffer@CommaOri t_{0}\ziffer@CommaOri\ldots\ziffer@CommaOri t_{2}$.
@@
We assess model performance on three different interaction types: quotes, replies, reposts and for three different engagement types: likes, replies, and reposts. As for baselines in comparison, we included DeepSets (zaheer2017deep), MLPs, and MPNNs. Note that the first two are connectivity-agnostic; unlike MLPs, DeepSets access, however, global feature information; see more in Section˜B.3.
@@ Results
Table˜3 shows a clear, consistent ordering from DeepSets to MLPs to MPNNs, across all interaction types in terms of lower MAE, higher $R^{2}$, and Spearman ($\sigma$). These results suggest that local, directed relational structures are informative for predicting engagements over a short time horizon, as both DeepSets and MLPs, which ignore graph structure, lag behind MPNNs. However, it has to be noted that the global feature information accessed by DeepSets appears detrimental. We also note that the absolute values of $R^{2}$ and $\sigma$ remain relatively low, indicating that the proposed tasks are far from being solved and leaving room for meaningful gains through more advanced, task-specific MPNNs.
@@ 3.2 Hardware design
Figure 3: Transformation of a Boolean formula into an And–Inverter Graph (AIG). The left side shows the logic network with explicit gates and inversions; the right side shows the corresponding AIG representation using only AND nodes (teal) and negated edges (orange).
@@ 3.2.1 Chip design: Learning to generate small circuits
Recently, machine learning has begun to influence key stages of chip design, including macro placement (alphachip) and floor planning (mallappa2024floorset). In this work, we propose focusing these efforts on logic synthesis, framing it as a conditional graph generation problem: generating a circuit that correctly implements a given Boolean function while minimizing the number of gates. The goal is thus to produce circuits that satisfy a prescribed truth table exactly while optimizing structural efficiency.
@@
Formally, let $f\colon\{0\ziffer@CommaMy 1\}^{n}\to\{0\ziffer@CommaMy 1\}^{m}$ be a Boolean function with $n$ inputs and $m$ outputs. Its behavior is fully specified by a truth table $\mathbold{T}^{f}\in\{0\ziffer@CommaMy 1\}^{m\times 2^{n}}$, where each column $\mathbold{T}^{f}_{\mathbold{x}}$ corresponds to an input $\mathbold{x}\in\{0\ziffer@CommaMy 1\}^{n}$ and satisfies $f(\mathbold{x})=\mathbold{T}^{f}_{\mathbold{x}}$. A logic circuit $C\coloneqq(V\ziffer@CommaOri E)$ is represented as a directed acyclic graph (DAG) of in-degree at most two. The node set is partitioned into input nodes $V^{\text{in}}\coloneqq\{v^{\text{in}}_{1}\ziffer@CommaOri\ldots\ziffer@CommaOri v^{\text{in}}_{n}\}$, gate nodes $V^{\text{gate}}\coloneqq\{v^{\text{gate}}_{1}\ziffer@CommaOri\ldots\ziffer@CommaOri v^{\text{gate}}_{k}\}$, each computing a Boolean function of arity at most two (e.g., And), and output nodes $V^{\text{out}}\coloneqq\{v^{\text{out}}_{1}\ziffer@CommaOri\ldots\ziffer@CommaOri v^{\text{out}}_{m}\}$, each of in-degree one. The edge set $E\subseteq V\times V\times\{0\ziffer@CommaMy 1\}$ specifies connections, where a label of $1$ denotes signal inversion (Not) and 0 denotes direct transmission. For an input $\mathbold{x}\in\{0\ziffer@CommaMy 1\}^{n}$, values are assigned to input nodes, propagated through gate nodes (applying inversion where specified), and collected at output nodes to yield the circuit output $C(\mathbold{x})\in\{0\ziffer@CommaMy 1\}^{m}$. Because the combination of And and Not is functionally complete, any Boolean function can be represented using only these two operations. We therefore adopt the standard representation of and-inverter graphs (AIGs), where circuits are DAGs of And gates connected by edges that may invert their signals (mishchenko2005fraigs). AIGs are widely used in industrial logic synthesis and provide a compact, canonical representation of circuits. Figure˜3 provides an overview of how the AIG is constructed from a Boolean formula and how it compares to traditional logic gates.
@@ Results
Figure 4: Transformation of an electronic circuit into a graph.
@@ 3.2.2 Electronic circuits: Predicting voltage conversion ratio
Analog circuit design remains critical but labor-intensive in EDA (Zhang2019CircuitGNNGN; Zhang2020GRANNITEGN; zhao2020automated). Unlike mature digital flows, analog design still relies on expert intuition due to sensitivity to device variations and complex constraints. This is especially true for custom power-converter design, where topology generation and parameter tuning must balance conflicting metrics. Even with a modest number of components, hundreds of thousands of topologies are possible; yet, new converter families are mostly handcrafted. Automating this process could accelerate design cycles and expand the exploration of novel architectures. Learning-based methods offer promise as surrogate models that predict performance (Fan2021FromST; fan2024graph), enabling faster search over vast spaces. Power converters’ natural graph structure makes them an ideal benchmark, highlighting robustness to structural sensitivity and the ability to cope with extreme class imbalance. Figure˜4 provides an overview of how an electronic circuit is transformed into a graph.
@@ 3.3 Reasoning and optimization
Figure 5: Overview of the three graph types generated from a given SAT formula.
@@ 3.3.1 SAT solving: Algorithm selection and performance prediction
In GraphBench, we introduce, to the best of our knowledge, the largest algorithm-selection benchmark for SAT solvers, covering eleven solvers and more than 100 000 instances. Our benchmark supports multiple graph representations of SAT formulae and also provides the SATzilla 2024 feature set (ShaHoo24), enabling their combination with MPNNs. Unlike existing GNN benchmarks for SAT (LiEtAl24), which primarily focus on satisfiability prediction or assignment generation, our benchmark is designed for practical downstream tasks in SAT solving—specifically, algorithm selection and performance prediction. Figure˜5 provides an overview of how the three graph types are constructed from a given Boolean formula.
@@
We created the largest algorithm-selection scenario for SAT solvers to date. We collected all SAT Competition instances via GBD (IseJab24) and added all instances from the AClib benchmark for algorithm configuration (HutEtAl14b). Together, these sources include SAT formulae from diverse real-world applications and synthetically generated instances. To further enrich the dataset, we generated additional formulae using three well-established SAT instance generators, i.e.,
@@
We provide three datasets, small, which includes only small formulae with up to 3 000 variables and 15 000 clauses, medium, which consists of the small formulae with additional medium size formulae with size of up to 20 000 variables and 80 000 clauses and large, which includes all formulae. This way, we provide datasets with varying hardness levels, as the size of the formula can pose hurdles for training time and GPU memory. While the large and medium datasets are unlikely to be accessible on current hardware, we provide them to keep the benchmark relevant in the long term.
@@
All SAT solving tasks in the benchmark are graph-level. For the graph transformer baselines, we employ an encoder-processor-decoder pipeline, as described in Appendix˜B, with no edge or node encoding. While we considered GT for the tasks, we encountered high computational costs in calculating the positional embeddings (more than a CPU day) and therefore excluded them from the evaluation. Furthermore, when training on medium- and large-sized graphs, we encountered out-of-memory errors even on an H100 with mixed-precision training. Consequently, we only consider small formulae for MPNNs. Similarly, we observed out-of-memory with the clause graphs, and therefore excluded them from the evaluation.
@@ Results
We start by presenting the results for the performance prediction task. We present the RMSE of the log-scaled computation time for the solvers from SAT Competition 2024 (Kissat, BreakIDKissat, KissatMABDC, AMSAT) for the small formulae in table Table˜8. The table shows that traditional empirical performance predictors (random forest and XGBoost) that use hand-crafted features outperform MPNNs by a high margin. Furthermore, we observe significant differences between the various graph types. While usage of LCG is common as an input for MPNNs in the context of SAT solving (SelEtAl19; TonGro25), the smaller VCG and VG perform better, which shows that the choice of input graph can have a strong influence on the results. We furthermore present the results of random forest and XGBoost on medium-size and large formulae in Tables˜9 and 9, such that they can be compared against graph machine learning solutions in the future. Overall, we observe that the random forest is the best-performing EPM for SAT solvers. For the algorithm selection task, we report the closed gap metric of different selectors in Table˜10. The VG representation allows for better performance than Both VCG and LCG, although using GIN across all three graph representation approaches, underperform pairwise regression with hand-crafted features. Furthermore, VG also outperforms multi-class classification and achieves a slightly smaller gap closure than pairwise classification. In contrast, using LCG leads to worse performance than the single best solver, indicating poor generalization to unseen instances. In their work, ZhaEtAl24 claimed that their MPNN-based method outperforms traditional approaches. However, there are several differences: the instance and solver composition differ, they used a SAT-specific MPNN layer, whereas we employ a general-purpose GIN, and our baselines use the latest SATzilla 2024 features, which have shown significantly better results than the previous version (ShaHoo24). Since the authors did not publish their code, we were unable to include their method as a baseline in our study. Among the three baselines based on hand-crafted features, we observe that multi-class classification underperforms both pairwise classification and regression, leaving a negative gap on the small and medium-sized datasets. This highlights that algorithm selection is not a simple multi-class classification problem. In that formulation, the model loses valuable information about the performance of algorithms other than the best one. Finally, we find more gaps closed for the large dataset across all three baselines, as it includes harder instances. As times increase, the benefit of algorithm selection grows, since feature computation accounts for a smaller fraction of total computation time. Moreover, for harder instances, multiple solvers frequently time out, so choosing a solver that successfully solves the instance becomes especially valuable, given the additional penalty imposed by the PAR score.
@@
Figure 6: Down-scaled sample graphs and solutions across all graph generators and all performed tasks for the Combinatorial Optimization datasets with our chosen parameters.